### v1

# Cluster analysis of earthquake hypocenters in Azerbaijan and surrounding territories

Abstract.The research focuses on seismic events that occurred in Azerbaijan and adjacent territories, regions known for strong seismic activity. We analyze a catalog of recorded earthquakes between 2010 and 2023, extracting the locations of the earthquake hypocenters for study purposes. Using statistical methods and cluster analysis tools, we developed a procedure for partitioning hypocenter clusters. The procedure begins with estimates of the Morisita Index, which is suitable for preliminary assessments of the statistical properties of hypocenter sets. Analysis of the Morisita Index indicates that the spatial distribution of hypocenters is heterogeneous, containing denser domains referred to as clusters. The next stage involves identifying spatial clusters using the DBSCAN and HDBSCAN algorithms. Due to the strong dependence of results on the algorithm’s parameters, we selected several partitions with 5–8 clusters that provided maximal or near-maximal Silhouette Index values. The final stage assesses the similarity of the resulting partitions, using the Adjusted Rand Index to identify partitions with a specified degree of similarity. The final set of partitions was compared to the fault network of the region. Based on the selected partition, the earthquake depth distributions were studied. Specifically, approximate probability density functions were constructed in the form of mixtures of normal distributions, leading to the identification of several bimodal distributions.
Keyword:Morisita index, Azerbaijan seismicity, focal depth distribution
Introduction
1 Introduction
The Caucasus is a region where the Arabian and Eurasian plates collide intensely. As highlighted by [1, 2, 3], this area is highly tectonically active, marked by the formation of seismogenic faults, fold-and-thrust belts, and depressions that are indicative of ongoing orogenic processes.
This tectonic activity gives rise to seismically active zones, characterized by a significant number of seismic events, including strong earthquakes with magnitudes reaching up to $M=7$ [3]. Despite the high seismic hazard in these regions, they are often densely populated, host critical infrastructure, and support economic activities, including the extraction of mineral and agricultural raw materials.
To comprehend the evolutionary scenarios of this geosystem, secure the sustainable functioning of the regional community, and develop strategies to mitigate seismic hazards, comprehensive studies of the Caucasus, particularly Azerbaijan, are consistently carried out.
In particular, these studies encompass the geology and geodynamics of the Caucasus [3, 4], magmatism and heat flow, active tectonics, and tectonic stresses (e.g., [3, 1, 4, 2]). They also address seismicity, gravity, and density models. Special attention is given to the exploration of active faults [5], including the Kur Fault, a major thrust structure within the Kura fold-and-thrust belt in Azerbaijan [6]. Notably, this region is not only seismically active but also home to critical infrastructure, such as the large Mingachevir and Shamkir hydroelectric water reservoirs and plants (e.g., [6, 7, 8]).
In addition to geological, seismological, and other applied research, theoretical studies of earthquake sequences utilizing statistical methods have been developed. The statistical analysis of Azerbaijan’s seismicity, observed from 2003 to 2016, was conducted by [9]. This analysis involved evaluating the frequency-magnitude distribution of earthquakes and fitting it to the Gutenberg-Richter law, constructing global and local coefficients of variation, extracting scaling exponents from magnitude time series, assessing the correlation dimension for the spatial distribution of epicenters, and examining other related parameters. However, the phenomenon of earthquake clustering remains insufficiently explored, despite its identification during analyses of the spatial distribution of earthquake epicenters (e.g., [9, 1, 2]).
It should be noted that recent advances in cluster analysis [10] enable more effective studies of this issue by mitigating the impact of large uncertainties in data, non-quantitative and subjective analyses, and the misinterpretation of significant amounts of inaccurate information [11, 12].
Numerous studies have been dedicated to earthquake clustering (e.g., [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]). This highlights the immense potential of cluster analysis and machine learning techniques across various scientific disciplines, particularly in seismology.
This research is based on the studies outlined in [25, 8]. It focuses on earthquakes that occurred in Azerbaijan and adjacent areas from 2010 to 2023, using an original procedure to cluster earthquake catalogs. Initially, the presence of formal clusters within an earthquake sequence is roughly estimated using the Morisita Index (MI), which serves as an indicator of grouped data. Following this, the catalog is divided into clusters using the DBSCAN [26, 22] and HDBSCAN [27] algorithms, with the results monitored through the Silhouette Index (SI). Partitions providing the maximal SI or values close to it are then compared for similarity using the Adjusted Rand Index (ARI) [28, 29, 30, 31], which helps analyze and reduce the number of partitions by eliminating those with high similarity. Additionally, the resulting partitions are compared with the regional fault network. This procedure enables a reliable identification of clusters representing earthquake epicenters or foci. Furthermore, the depth distributions of earthquake clusters are examined, and these analyses contribute to the development of a systematic understanding of seismic activity.
2 The earthquake data selection
We examine earthquakes occurring within the rectangular region bounded by latitudes 38 to 42 N and longitudes 45 to 50 E. From the catalog [32], a total of 6,194 earthquakes with magnitudes $M\geq 2$ and depths exceeding 0.1 km were identified. A comparable dataset of Azerbaijani earthquakes has also been analyzed by [9]. Each seismic event is defined by its latitude, longitude, depth, time, and various magnitude types.
As a rule, information extracted from earthquake catalogs requires pre-processing [23], which typically involves the removal of non-tectonic events [24, 33] and those assigned characteristics by the catalog operator [33]. An essential step in catalog processing is the verification of its completeness, which includes the estimation of the completeness magnitude $M_{c}$ [9, 23].
This study does not involve filtering the catalog data. Instead, we directly proceed to estimate the completeness magnitude $M_{c}$, which represents the minimum magnitude above which all earthquakes are reliably recorded in the catalog. To determine $M_{c}$, we applied the MAXC method [34, 9]. Based on our previous analysis [8], a completeness magnitude of $M_{c}=2$ is appropriate, allowing us to include all earthquakes with $M\geq M_{c}$ in the analysis.
To provide an overview of the spatial distribution of the selected earthquakes, we plot the earthquake epicenters in Fig. 1a as filled circles, with color indicating the depth of each event. Most earthquakes are relatively shallow, while deeper seismic events are predominantly concentrated near the coast.
(a)                                                                       (b)
Figure 1: a: Distribution of earthquake epicenters across Azerbaijan and surrounding regions. Earthquake depth is represented by color according to the accompanying colorbar. b: Normalized locations of the earthquake epicenters corresponding to panel (b).
3 Morisita index application
Since our focus is on identifying earthquake groupings, it is appropriate to begin with a coarse criterion for clustering. To this end, we apply an approach that estimates the dispersion of spatially distributed points. Specifically, we use the box-counting method, in which the study area is divided into small cells—typically square grids—and the number of points within each cell is counted. This provides a basis for introducing the Morisita Index [35, 36, 37], a quantitative measure of spatial clustering.
$I_{\delta}=Q\frac{\sum_{i=1}^{Q}n_{i}(n_{i}-1)}{N(N-1)},$ (1)
where $Q$ is the number of grid cells (boxes) covering the study area, $n_{i}$ denotes the number of points within the $i$th cell, and $N$ is the total number of points in the area.
The Morisita Index can be formally extended to multi-dimensional data. However, the computational complexity increases significantly with dimensionality. Therefore, it is practical to perform preliminary calculations using two-dimensional data representing earthquake epicenter locations, and subsequently compare the results with those obtained from three-dimensional data (hypocenter coordinates), where applicable.
To simplify the evaluation of $I_{\delta}$, we transform the coordinates of earthquake hypocenters $(P,R,H)$ into the unit cube using the following mapping:
$x=\frac{P-P_{\min}}{P_{\max}-P_{\min}},\quad y=\frac{R-R_{\min}}{R_{\max}-R_{\min}},\quad z=\frac{H-H_{\min}}{H_{\max}-H_{\min}},$ (2)
where $\{P_{\min},P_{\max}\}=\{41.700;52.128\}$, $\{R_{\min},R_{\max}\}=\{37.177;42.482\}$, $\{H_{\min}$, $H_{\max}\}=\{0;112.1\}$. Then each side of the cube is divided into $n$ parts of the length $\delta=1/n$.
We begin the MI calculation with the case of 2D data, which, after the transformation (2), is distributed over the unit square, as shown in Fig.1b. The unit square is subdivided into $Q=n^{2}$ quadrats. The resulting Morisita index plot is presented in Fig. 2a (the lower broken line), illustrating the trend as the number $n$ increases from 1 to 100, or equivalently, as the mesh size $\delta$ decreases from 1 to 1/100.
The graph is bounded on the right by the value $I_{\delta}=1$ at $\delta=1$ (or $n=1$). In particular, local minima (”pits”) are observed at $n=m_{1}=8$, $n=m_{2}=12$, etc. To investigate the nature of these features, auxiliary studies were conducted.
To better understand the emergence of ”pits”, we compare the positions of two meshes obtained at $n=m_{2}=12$ and $n=11$, demonstrating their relation to cluster capture. The zoomed-in section of the unit square from Fig. 1b, which contains both meshes and a cluster, is shown in Fig. 2b.
Consider the partial quadrat formed by the less dense mesh, assuming that it contains $r$ cluster points, with points outside the square excluded from consideration. The denser mesh subdivides this quadrat into four smaller quadrats, each containing fewer cluster points—denoted as $r_{1,2,3,4}$—such that $r=\sum r_{i}$. Then the expression (1) for $I_{\delta}^{11}$ incorporates the term $r(r-1)$, while $I_{\delta}^{12}$ contains $\sum r_{i}(r_{i}-1)$. Let us examine their difference $r(r-1)-\sum r_{i}(r_{i}-1)=r^{2}-\sum r_{i}^{2}>0$. From this it follows that $I_{\delta}^{11}>I_{\delta}^{12}$. The above considerations offer a qualitative framework for understanding how the Morisita Index effectively distinguishes a sufficiently contrasting cluster from a background of sparsely distributed points.
(a)                                                                       (b)
Figure 2: a: Morisita Index $I_{\delta}$ vs. $\delta$. Here $m_{1}=8$ and $m_{2}=12$. Inset shows the sequence $L_{n}=\ln I_{1/n}/\ln n$, defined in (3), approaching its constant value (green). The plots are evaluated for 2D (black) and 3D (red) data, respectively. b: The part of Fig.1b with meshes corresponding to $\delta=1/m_{2}=1/12$ (red) and $\delta=1/11$ (green).
Transforming the locations of earthquake hypocenters using relation (2), we perform similar MI calculations for the 3D data, which allows us to construct the upper broken line in Fig. 2a. Unfortunately, the calculation of $I_{\delta}$ was terminated after $n=30$ due to the high computational demands. Nevertheless, comparing the shapes of the $I_{\delta}$ curves obtained for the 3D and 2D data reveals a clear similarity. Moreover, the ”pits” become more pronounced, although their positions remain unchanged.ronounced, while their positions remain unchanged.
Thus, since the MI exceeds 1, we can conclude that earthquakes are distributed non-uniformly and tend to form clusters. Before proceeding to cluster identification, let us consider another feature of MI for the sake of completeness.
Moreover, MI analysis provides additional insight into the fractal nature of datasets [23]. The method of MI calculation—namely, the box-counting approach—is also commonly used to assess the dimensions of fractal objects. Therefore, it is not surprising that MI is related to the fractal dimensions of self-similar point patterns. As shown by [36], the relationship between MI and the fractal dimension is established through an asymptotic power-law dependence, $I_{\delta}\propto\delta^{p}$ ($p<0$). According to [36], the exponent $p$ is related to the second-order Morisita slope (also referred to as the codimension of order 2) by $p=-S_{2}$. Furthermore, the relation $D_{2}-E=p$ holds, where $D_{2}$ is the second-order Rényi generalized dimension (or correlation dimension) and $E$ is the dimension of the Euclidean space in which the points are embedded. In addition, the following relation is valid (with $\delta=1/n$):
$\lim_{n\to\infty}L_{n}=\lim_{n\to\infty}\frac{\ln I_{1/n}}{\ln n}=p=const.$ (3)
To verify property (3), we evaluate the sequence $L_{n}$ for sufficiently large $n$, specifically for $n>10$. The resulting dependencies are shown in the inset of Fig. 2a. It can be observed that $L_{n}$ tends toward a constant value, with $p\approx-0.68$ for the 2D data and $p\approx-1.27$ for the 3D data. This indicates that the spatial distribution of earthquakes exhibits fractal properties.
It is worth noting that [9] calculated the correlation dimension for spatially distributed epicenters to be approximately $D_{c}\approx 1.38$. Our evaluations yield a close estimate of the fractal dimension, with $D_{2}=2+p=1.32$. For the 3D case, the corresponding dimension is $D_{3}=3+p=1.73$.
4 DBSCAN and HDBSCAN application
To perform the cluster analysis, we use the well-known DBSCAN algorithm and its improved version, HDBSCAN, implemented via the sklearn.cluster module in Python [38]. Some preliminary results concerning the application of DBSCAN to this catalog were previously presented in [25]. Here, we summarize only the key points from [25].
To improve the results, we preprocess the data by scaling the earthquake depths according to the relation $Z=z/|z|$, where the Euclidean norm is defined as $|z|=\sqrt{\sum_{j}z_{j}^{2}}$.
Next, the DBSCAN algorithm is applied to the catalog. Recall that DBSCAN ([26, 16, 22]) requires two intrinsic parameters: $\varepsilon$ and min_samples. The former defines the $\varepsilon$-neighborhood of each point, specifying the maximum distance between any two points within the neighborhood. The latter sets the minimum number of points required within the $\varepsilon$-neighborhood to form a cluster. To estimate appropriate ranges for these parameters, we apply the approach proposed by [26], with an adaptation detailed in Appendix A.
Thus, the parameter $\varepsilon$ is selected from the range $\varepsilon\in[0.2,0.6]$ with a step size of 0.05, and min_samples is chosen from the range $[80,160]$ with a step size of 5. Two types of distance metrics are used—”euclidean” and ”manhattan”—resulting in a total of 306 cases.
The result of DBSCAN is a partition consisting of clusters along with a set of unclassified points (noise points). It is important to note that different combinations of the parameters $\varepsilon$ and min_samples can yield partitions with the same number of clusters.s.
The quality of point classification is assessed with the help of Silhouette Index [39], the set of which is shown in Fig.3a. To consider the influence of the parameter $\varepsilon$ to clustering, we associate point positions with a colorbar such that the lighter points correspond to higher values of $\varepsilon$. From Fig.3a it follows that increasing $\varepsilon$ leads to the reducing the number of clusters. A similar effect is observed with increasing min_samples.
Table 1: The DBSCAN’s parameters corresponding to the maximum SIs depicted in Fig.3a.
Number of clusters 4 5 6 7 8
$\varepsilon$ 0.5 0.4 0.45 0.45 0.4
min_samples 140 140 130 135 110
(a)                                                                       (b)
Figure 3: a: Silhouette Index vs. number of clusters. b: The 6-cluster partition corresponding to the maximal Silhouette Index depicted in the left panel by “$\times$”. The noise points are removed. Filled bullets mark earthquakes with $M>5.5$. The results obtained via DBSCAN.
(a)                                                                       (b)
Figure 4: a: Silhouette Index vs. number of clusters. b: The 6-cluster partition with noise points removed, corresponding to the maximal Silhouette Index, depicted in the left panel by the symbol “$\times$”. Filled bullets mark earthquakes with $M>5.5$. The results obtained via HDBSCAN.
Regarding the application of DBSCAN, it is worth making some remarks about the selection of the intrinsic parameter values. As shown in Fig.2a, MI profiles exhibit ”pits” that correspond to crucial changes in the structure of the dataset partition. In other words, MI can detect the approach of a cluster. This suggests that we can approximate the spatial scale for the mesh containing a cluster. For example, the first local pit is observed at $\delta=1/8$, and the quadrat containing the cluster has a side length of $\delta=1/7$. In the initial coordinates, the corresponding quadrat can have a side length of $\min{\delta(P_{\max}-P_{\min});\delta(R_{\max}-R_{\min})}=0.758$. The parameter $\varepsilon$ for DBSCAN defines a region with a side length of approximately $2\varepsilon$, which falls within the interval $(0.8;1.0)$ (Table 1) for cases where $4\leqslant N\leqslant 8$, and the SI approaches its maximum. There is also the case where $\varepsilon=0.3$ ([25]), which corresponds to a 7-cluster partition and lies close to SImax. Thus, the observed correlation between MI properties and DBSCAN parameters may not be accidental and can assist in applying the DBSCAN algorithm more effectively.
[…]
5 Adjusted Rand Index analysis
To quantify the differences between the partitions, we use the Adjusted Rand Index (ARI), which measures the similarity between two clusterings. Generally, these two partitions may consist of different numbers of subsets; however, in this study, we compare partitions containing the same number of clusters. It is important to note that the ARI ranges from $-1$ (indicating particularly discordant partitions) to 1 (indicating identical partitions). For the ARI evaluation, we use the adjusted_rand_score function from the sklearn.metrics module [40]. This approach helps us reduce the number of partitions that need to be considered as different.
In the following, we select partitions with a fixed number of clusters, $N$, which are compactly located and possess a sufficiently large SI. We begin with DBSCAN partitions and consider those with $N=5$ clusters, selecting the subset enclosed in the rectangle shown in Fig.3a. Similarly, partitions for $N=6$ and $N=7$ are also selected. The results of the ARI evaluation for pairs of clusterings obtained for $N=5$, $N=6$, and $N=7$ are presented in matrix form in Figs.5a, 5b, and 5c, respectively. The row and column labels correspond to the partitions for which the ARI is evaluated, listed in the order of their appearance within the specified rectangle in Fig.3a. The evaluated ARI values are placed in the corresponding cells and highlighted by color. ARIs for identical partitions are equal to 1 and are depicted in dark colors, while other combinations yield smaller ARIs, represented in lighter shades. We observe that the ARI varies within the range $[0.86;1]$ for the diagram in Fig.5a, within $[0.92;1]$ for the diagram in Fig.5b, and within $[0.72;1]$ for the diagram in Fig. 5c.
(a)                                                                       (b)
Figure 5: Adjusted Rand Index evaluation for 5- (a), 6- (b), and 7-cluster partition (c). The axis labels correspond to the selected partitions in Fig. 3a, numbered in descending order of SI, in particuler, the partition marked by “$\times$” correponds to the label 12 in the panel (b). The results obtained via DBSCAN.
(a)                                                                       (b)
Figure 6: Adjusted Rand Index evaluation for 5-cluster partition (a) and 6-cluster partition (b). The axis labels correspond to the selected partitions in Fig. 4a, numbered in descending order of SI, in particular, the partition marked by “$\times$” correponds to the label 0 in the panel (b). The results obtained via HDBSCAN.
Similarly, the ARI matrices are calculated for HDBSCAN partitions consisting of $N=5$ clusters and $N=6$ clusters, which are contained within the rectangles shown in Fig.4a. According to Fig.6a, the ARI fluctuates within the interval $[0.79;1]$ for the 5-cluster partitions. For the 6-cluster partitions, as shown in Fig.6b, the ARI takes values from the set ${0.86,0.87,0.88}$.
Such high ARI values indicate a strong similarity between the selected partitions. The closer the ARI is to 1, the fewer differences can be distinguished between the partitions. For instance, when comparing two partitions with the same number of clusters, the differences may manifest in the location of individual points, or small portions of one cluster may overlap with another cluster.
Thus, the ARI evaluation allows us to categorize the selected partitions into two classes: those that are similar to each other and those that are significantly different. In other words, a threshold ARI value (which we have determined) can be introduced to distinguish between partitions that are most similar from an informational perspective. Therefore, to represent the similar partitions, we can select any one of them.
Further reduction in the number of partitions requires additional information about the earthquakes. In particular, we compare the clusters with the structure of the fault network. For the 6-cluster partitions obtained using both DBSCAN and HDBSCAN, we select the partitions that provide the maximum SI values. The locations of the clusters and the fault network are shown in Fig.3b and Fig.4b. A good agreement is observed between the cluster locations and the zones with denser and more diverse fault subsystems.
The selected partition should not be considered the only possible one. Other 6-cluster partitions, which we compared using ARI, can also be used.
6 Analysis of the earthquake depth distributions
Using the evaluated clusters, various statistical properties of local seismicity can be estimated, including the depth distribution of earthquake hypocenters.
(a)                                                                       (b)
Refer to caption   Refer to caption
(c)                                                                       (d)
Refer to caption   Refer to caption
(e)                                                                       (f)
Figure 7: Histograms of the earthquake depth calculated for total number of earthquakes and earthquakes belonging to the specified cluster. Solid and dashed curves stand for the approximating PDFs. Clusters are taken from Fig.4b.
Note that studies of earthquake depth distributions are crucial for understanding the structure of the lithosphere, its rheology, strength, and stress state [41, 42], as well as clarifying the temperature distribution in the Earth’s upper layers [42], assessing the impact of surface activity on local seismicity [43], and more.
As a rule, most earthquakes occur within a single crustal layer [42], known as the seismogenic crust, which lies above the Moho and is approximately 10 to 40 km thick [41]. Typical histograms of local earthquakes [44, 45, 46, 42] are characterized by unimodal distributions with a pronounced peak at depths ranging from 10 to 20 km. A similar pattern has also been observed in Azerbaijani seismicity, particularly in the Kura fold-and-thrust belt and the surrounding area, as noted by [6]. However, the depth distribution for the total number of earthquakes worldwide, occurring at depths up to 700 km [47], has been found to be bimodal. Furthermore, shallow seismicity observed in subduction zones, such as the Eastern Aleutians, Kermadec Islands, Solomon Islands, and others [44], as well as in the Charlevoix seismic zone (Canada)[48] and western Japan [49], also exhibits bimodal depth distributions.
Thus, let us construct the earthquake depth histograms for the clusters shown in Fig.4b. For comparison, the histogram of all earthquakes is also depicted in Fig.7. All histograms exhibit a maximum around 10 km, with the upper cutoff depth at 60 km and lower cutoff depths for the clusters. This is in good agreement with other seismic research (e.g., [6]) and the assessment of Moho depth in Azerbaijan, which, according to [4], varies from approximately 40 km in the Kura basin and northern foreland basin to over 50 km in the eastern Greater Caucasus.
As shown by [44], to fit the histograms (see details in Appendix B), the probability density functions (PDFs) of a mixture of normal distributions are used. The application of this PDF fitting procedure revealed that attempts to approximate the histograms in Figs.7c and 7d using a mixture of normal distributions were unsuccessful. Instead, a single normal distribution was found to be more appropriate. Interestingly, the histograms in Figs.7e and 7f exhibit a bimodal nature. Numerical verification shows that the PDF for the histogram of the total number of earthquake depths has two modes, located at approximately 10 km (intense) and 30 km (almost indistinguishable). The latter peak becomes more pronounced in the cluster histograms (Figs.7e, f). In particular, the deeper peak in cluster 6 (Fig.7f) is located at a depth of around 35 km. The studies by [50] on seismicity in the Ismayilly area, which is covered by cluster 6, also confirm the accumulation of hypocenters at a depth of about 40 km.
As indicated by [49], the two modes observed in the depth histogram for Japanese earthquakes reflect distinct sources: the shallow peak (around 5 km) is associated with crustal earthquakes and anthropogenic events, while the deeper peak (around 35 km) is attributed to seismic events near the plate boundary and deep crustal earthquakes. Other factors contributing to the formation of these two modes may be related to the presence of complex tectonic structures, as discussed by [2] and [50]. Overall, this issue requires a comprehensive analysis of geological and geophysical data for the region.
7 Concluding remarks
In this research, the seismicity of Azerbaijan was studied, with a particular focus on developing a formal strategy for clustering the spatial distribution of earthquake hypocenters. Preliminary evidence of the presence of clusters was obtained by analyzing the Morisita Index, which reliably indicates the potential for cluster selection. Additionally, the Morisita Index evaluation provided an estimate for the fractal dimension of the distributions of both earthquake epicenters and hypocenters.
Using the DBSCAN and HDBSCAN algorithms, we evaluated partitions of the earthquake catalog. Since there are no strict restrictions on the intrinsic parameters of the algorithms, we obtained a variety of partitions containing between 2 and 20 clusters. To assess the quality of clustering and select the appropriate partitions for further study, the Silhouette Index was calculated. Partitions providing the maximum Silhouette Index, or those close to it, were selected. The next step in reducing the number of partitions involved using the Adjusted Rand Index. This index was calculated for pairs of partitions with the same number of clusters, allowing us to identify similar partitions. As a result of these clustering stages, we selected partitions containing 5-8 clusters, which are the most informative from a data perspective.
Since we performed a formal analysis of the dataset, additional earthquake-related information should be used to make the final partition selection. In this study, we compared the partitions with the regional fault network. This comparison led to the selection of a partition containing 6 clusters, which aligns well with the regional fault network.
Using the selected partition, we studied the earthquake depth distributions by constructing probability density functions (PDFs) for each cluster. To approximate the PDFs, we employed a mixture of normal distributions. It was found that the clusters located along the sea coast are characterized by bimodal distributions.
Data availability
The earthquake catalog used in the current study is available at http://www.isc.ac.uk/iscbulletin/search/. The fault catalog is available at http://neotec.ginras.ru/index/english/database_eng.html. Python source code implementing the cluster analysis of the earthquake catalog is available at https://github.com/SkurativskaKateryna/IJ_NLM_Earthquake_clusterization.git.
Appendix A Estimation of DBSCAN parameters
The selection of the parameters $\varepsilon$ and $Min\_samples$ for the DBSCAN algorithm is based on the analysis of the sorted $k$-distance graph, as discussed by [26] and [15], which uses the sorted distance matrix. We plot the $k$-distance graphs for $k=4$, $k=70$, and $k=150$ (Fig. 8). The appropriate value of $\varepsilon$ corresponds to the formation of the ”knee” on these graphs. Therefore, we choose $\varepsilon$ from the range $[0.2;0.6]$ and $Min\_samples$ from the range $[25;150]$. It is important to note that this procedure should still be considered an approximate method for parameter estimation, as noted by [16] and [51].
Figure 8: The sorted $k$-distance graphs.
Appendix B Fitting the earthquake depth histograms by a mixture of normal distributions
The histograms shown in Fig.7 are approximated by a mixture of normal distributions. To do this, we used the Mathematica 12.0 in-built procedure EstimatedDistribution[data, MixtureDistribution[{a, b}, NormalDistribution[$\mu_{1}$, $\sigma_{1}$], NormalDistribution[$\mu_{2}$, $\sigma_{2}$]]], where $data$ stands for the array of depths, while the parameters $a$, $b$ (it is obviously that $a+b=1$, but we follow the Mathematica syntax), $\mu_{1,2}$, and $\sigma_{1,2}$ should be estimated. Then the resulting PDF coincides with a linear combination of the normal distributions $N(\mu_{1,2},\sigma_{1,2})$. The calculated values of parameters are presented in Table 2.
Table 2: The parameters for the mixture of normal distributions depicted in Fig.7.
$a$ $b$ $\mu_{1}$ $\sigma_{1}$ $\mu_{2}$ $\sigma_{2}$
Total depths 0.674 0.326 9.945 4.842 30.147 15.137
Fig.7(a) 0.644 0.356 11.148 5.011 24.235 10.126
Fig.7(b) 0.937 0.063 8.140 3.749 21.719 9.272
Fig.7(c) 9.892 4.291
Fig.7(d) 9.614 4.986
Fig.7(e) 0.698 0.302 2.544 4.362 8.444 3.223
Fig.7(f) 0.732 0.268 9.166 4.849 34.938 10.027
It turned out that this procedure can fail when $b$ is close to zero. Then we applied EstimatedDistribution[data, NormalDistribution[$\mu_{1}$, $\sigma_{1}$].